OpenStudy (anonymous):

Prove the identity (sinx)(tanx cosx - cotx cosx) = 1- 2 cos^2x

zepdrix (zepdrix):

Rewrite \(\large \tan x\) and \(\large \cot x\) in terms of sines and cosines. Then cancel some stuff out. And simplify.\[\large \sin x\left(\tan x \cos x- \cot x \cos x\right) \qquad = \qquad \sin x\left(\frac{\sin x}{\cos x} \cos x- \frac{\cos x}{\sin x} \cos x\right)\]

OpenStudy (anonymous):

how come this stuff never shows up when people use the symbols? I cant read it :(

OpenStudy (anonymous):

i really appreciate your help, i just cant see anything

zepdrix (zepdrix):

Hmm what browser are you using? :( I think you'll have problems if you're on internet explorer.

OpenStudy (anonymous):

oh that must be it! i am on that, i will switch quick and see if it shows up! thank you so much!

zepdrix (zepdrix):

WHUT! Why would you EFER use Internet Explorer? +_+ Is it your first day on the interwebs? :O

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

bahahaha no, its just what ive always used. don't worry i get made fun of for using it all the time. and the computers at my school have that

zepdrix (zepdrix):

fair enough :3

OpenStudy (anonymous):

ok @zepdrix help! i went from (sinx)(tanx cosx - cotx cosx) to (sinx)(sinx/cosx * cosx - cox/sinx *cosx) to (sinx)(sinx - cos^2x/sinx) NOW WHAT?

zepdrix (zepdrix):

Hmm in the second set of brackets.. let's look at the first fraction there. It looks like we have a couple of cosines that will cancel out, don't we?

zepdrix (zepdrix):

Are my pretty graphics showing up now, or not so much? :o

OpenStudy (anonymous):

kind of, but it is like cut off :/

zepdrix (zepdrix):

Hehe yah, my bad.

zepdrix (zepdrix):

\[\large \sin x\left(\frac{\sin x}{\cancel{\cos x}} \cancel{\cos x}- \frac{\cos x}{\sin x} \cos x\right)\]So we have a nice cancellation here right?

OpenStudy (anonymous):

yep did that! then i multiplied cosx/sinx times cosx/1

zepdrix (zepdrix):

\[\large \sin x\left(\sin x- \frac{\cos^2 x}{\sin x}\right)\]Hmm, Ok cool. Next thing we'll want to do is `distribute` the \(\large \sin x\) to each term in the brackets.

OpenStudy (anonymous):

which is sin^2x - cos^2xsinx/sin^2x correct?

zepdrix (zepdrix):

Yep, see any other cancellations? c;

zepdrix (zepdrix):

Woops, how'd u get a sin^2x on the bottom of your second fraction? :O

zepdrix (zepdrix):

Remember, you were multiplying it by \(\large \dfrac{\sin x}{1}\)

OpenStudy (anonymous):

(sorry, question here, are your little symbols just supposed to be lines and fancy x's?)

OpenStudy (anonymous):

oh woops brain fart there. forgot that you dont multiply that on top and bottom - its over 1

zepdrix (zepdrix):

Lines and fancy x's?? XD lol The last fancy thing I sent should look like a fraction. It's just to format the math so it's a lot easier to read.

OpenStudy (anonymous):

ok so i cancelled and am now at sin^2x - cos^2x

OpenStudy (anonymous):

haha that is what i mean - lol fractions ok because that is what im seeing so its all good

OpenStudy (anonymous):

what do i do next?

zepdrix (zepdrix):

We want to remember back to our most basic Trig Identity \(\large \sin^2x+\cos^2x=1\). If we subtract cos^2x from each side, it gives us, \(\large \sin^2x=1-\cos^2x\). It appears we're trying to get an answer that only involves cosine. So maybe we can use this identity. Hmmm?